OpenStudy (anonymous):

What is the radius of a circle with the equation x2 + y2 – 14x + 10y = 250?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x^2 + y^2 – 14x + 10y = 250 }\) \(\normalsize\color{blue}{ x^2 – 14x+ y^2 + 10y = 250 }\) \(\normalsize\color{blue}{ x^2 – 14x\color{red}{+49} + y^2 + 10y\color{green}{+50} = 250\color{red}{+49} \color{green}{+50} }\) \(\normalsize\color{blue}{ x^2 – 14x-49+ y^2 + 10y +50= 349 }\) \(\normalsize\color{blue}{ (x-7)^2 + (y+5)^2= 349 }\)

OpenStudy (solomonzelman):

Compare \(\normalsize\color{black}{ (x-7)^2 + (y+5)^2= 349 }\) to \(\normalsize\color{black}{ (x-h)^2 + (y-k)^2= r^2 }\)

OpenStudy (bradely):

x^2 + y^2 – 14x + 10y - 250=0 Equation is in the form of x^2 + y^2+2gx +2fy+c =0 Formula for radius r =sqrt(g^2+f^2-c) 2g= -14 g=-7 2f=10 f=5 r=sqrt((-7)^2+(5)^2+250) =sqrt(49+25+250) =18 source: http://www.mathskey.com/question2answer/